# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ III-A Unified ISAC Model
  • UE-to-BS bistatic ISAC mode: For UE-to-BS bistatic ISAC mode, the ISAC-TX in Fig. 2 corresponds to a UE, while the Com-RX and Sen-RX are collocated at the BS, as shown in Fig. 3(b). Note that unlike the BS monostatic mode in Fig. 3(a), since only signal reception needs to be performed at the BS in Fig. 3(b), the Com-RX and Sen-RX may use the same set of antennas. However, synchronization is essential due to the use of different local oscillators (LO) and reference clocks by the ISAC BS and the UE. Furthermore, the ISAC signal $\mathbf{x}(t)$ in (1)-(3) is transmitted by the UE, and is denoted as $\mathbf{x}_{\mathrm{UE}}(t)\in\mathbb{C}^{M_{t}\times 1}$.
@@
  • BS-to-BS bistatic ISAC mode: For BS-to-BS bistatic ISAC mode, the ISAC-TX in Fig. 2 corresponds to a BS, while the Com-RX and Sen-RX occur at a UE and another BS, respectively, as shown in Fig. 3(c). Note that unlike UE-to-BS bistatic mode in Fig. 3(b), although BS bistatic mode also requires synchronization, it can be achieved through a wired connection between the two ISAC BSs. Furthermore, the ISAC signal $\mathbf{x}(t)$ in (1)-(3) is broadcast by the transmit BS, and is denoted as $\mathbf{x}_{\mathrm{BS}}(t)\in\mathbb{C}^{M_{t}\times 1}$.
@@
The corresponding ISAC-TX, Com-RX, and Sen-RX of the unified model in Fig. 2, as well as the required processing for the six ISAC modes, are summarized in Table II.
@@
In the following, we focus on ISAC systems implemented with OFDM waveforms, which is the dominant waveform in 4G and 5G communication systems, and is likely to continue to play a critical role for 6G. In OFDM-ISAC systems, target sensing can be performed based on pilots only or on the fully reconstructed transmitted OFDM data frame. The former approach can be easily implemented, but its sensing performance is limited by the availability of time-frequency resources allocated to pilots. The latter can fully exploit all of the time-frequency resources for target sensing using known (for monostatic sensing) or demodulated communication symbols (for bistatic sensing) for parameter estimation, at the expense of increased complexity. In this paper, we employ the OFDM frame-based approach for sensing tasks using all time-frequency resources allocated to both pilots and communication data.
@@ III-B MIMO-OFDM ISAC Signal Processing and Modeling
For MIMO-OFDM ISAC, the ISAC-TX transmits OFDM signals over $N$ subcarriers and $P$ OFDM symbols. The subcarrier spacing and the OFDM symbol duration without the CP are denoted by $\Delta f$ and $T$, respectively, where $T=\frac{1}{\Delta f}$. The system bandwidth is thus $B=N\Delta f$. Let $T_{\mathrm{cp}}$ denote the CP duration, so the OFDM symbol duration including CP is $T_{s}=T+T_{\mathrm{cp}}$. To avoid ISI for communication, the CP duration should be no smaller than the multi-path delay spread $\tau_{\mathrm{com}}$ of the communication channel $\mathbf{H}_{c}(t,\tau)$, since the Com-RX only begins receiving communication signals upon receiving its first path. However, for target sensing, the CP duration should be larger than the propagation delay $\tau_{\mathrm{sen}}$ corresponding to the ISI-free sensing distance [106], which is different from the case of communications since the Sen-RX begins to receive signals as soon as the ISAC-TX transmits, so that no nearby targets are missed. Thus, in ISAC systems, the CP should be set as $T_{\mathrm{cp}}\geq\tau_{\max}$, where $\tau_{\max}=\max\{\tau_{\mathrm{com}},\tau_{\mathrm{sen}}\}$. Thus, the OFDM signal $\mathbf{x}(t)\in\mathbb{C}^{M_{t}\times 1}$ transmitted by the $M_{t}$-antenna ISAC-TX array is given by
@@
×$\qquad\quad\times e^{j2\pi\upsilon_{k}t}\mathrm{rect}\left(\frac{t-\tau_{k}-pT_{s}}{T_{s}}\right)+\mathbf{z}(t),$ (5)
@@
The Sen-RX samples the sensing signal with sampling rate $B=N\Delta f$. To avoid inter-carrier interference (ICI) caused by large Doppler frequency shifts, the subcarrier spacing is typically assumed to be much larger than the Doppler frequency shift, i.e., $\Delta f\gg|\upsilon_{k}|,\forall k=1,2,\cdots,K$. Under this assumption, the $q$-th ($q=0,1,\dots,N-1$) sample of the $p$-th block can be expressed as
@@
×$\qquad\qquad\times e^{j2\pi\upsilon_{k}\left(q/B+pT_{s}+T_{\mathrm{cp}}\right)}+\mathbf{z}^{p}\left[q\right]$
@@
×$\qquad\qquad\qquad\qquad\times e^{j2\pi pT_{s}\upsilon_{k}}+\mathbf{z}^{p}[q],$
@@
Parameter estimation is one of the most important sensing tasks in ISAC systems, as it underpins functions like target localization, tracking, and imaging. Therefore, in the following, we will focus on the estimation of the location $\mathbf{q}_{k}$ (or the angle $\theta_{k}$ in the far-field case), the propagation delay $\tau_{k}$, and the Doppler frequency shift $\upsilon_{k}$ of the sensing targets. Generally speaking, the three parameters of each target can be estimated from the spatial, delay and Doppler domains. Inspired by traditional joint parameter estimation methods [107, 108], we propose four frameworks to estimate target parameters across the three domains. As shown in Fig. 4, depending on whether the parameters are estimated jointly or separately in each domain, the four frameworks are parallel one-domain, sequential one-domain, joint two-domain and joint three-domain parameter estimation. The parallel one-domain approach separately estimates the sets of target locations $\mathbb{S}_{\mathbf{q}}=\{\mathbf{q}_{1},\cdots,\mathbf{q}_{K}\}$ or angles $\mathbb{S}_{\theta}=\{{\theta}_{1},\cdots,{\theta}_{K}\}$, delays $\mathbb{S}_{\tau}=\{\tau_{1},\cdots,\tau_{K}\}$ and Doppler shifts $\mathbb{S}_{\upsilon}=\{\upsilon_{1},\cdots,\upsilon_{K}\}$, and then associates the individual parameter estimated into groups $(\mathbf{q}_{k}/\theta_{k},\tau_{k},\upsilon_{k})$ for each of the $K$ sensing targets. This approach has a relatively low computational complexity, but its sensing performance is limited by potential errors in the parameter grouping.
@@
On the other hand, the sequential one-domain approach performs spatial, delay and Doppler domain estimation sequentially. For example, spatial domain estimation can be applied first to determine the $K_{\hat{\mathbf{q}}}$ distinguishable locations/angles. When the spatial resolution is limited, it may not be possible to distinguish all $K$ targets in the spatial domain alone. Thus, in general, $K_{\hat{\mathbf{q}}}\leq K$. After location/angle estimation, beamforming can be applied to separate the $K_{\hat{\mathbf{q}}}$ signals in the spatial domain, denoted as $\mathbf{X}_{k_{\hat{\mathbf{q}}}},k_{\hat{\mathbf{q}}}=1,\cdots,K_{\hat{\mathbf{q}}}$. Subsequently, delay and Doppler domain processing can be applied. Note that the order in which the different types of parameters are estimated depends on the resolution and target separation in the three domains. The sequential one-domain estimation method does not require parameter grouping, but its complexity is greater than parallel one-domain processing.
@@
The details of each method will be discussed in Sections IV and V for the far-field and near-field cases, respectively. Note that the TX and RX steering vectors $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})$ and $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ are functions of target location $\mathbf{q}_{k}$, which in general depend on both the angle of departure (AoD)/AoA $\theta_{k}$ and distance $r_{k}$ between the target and the reference antenna. In the following, we present expressions for the RX steering vectors for the far-field and near-field scenarios. Similar results can be obtained for the TX array response.
@@ III-C Array Response Vector: From Far-Field to Near-Field
(b) Non-uniform spherical wave.
@@
When the BS array aperture is small, targets are typically located in the far-field region, and the conventional UPW model accurately characterizes the variations of signal phases across the array, as shown in Fig. 5(a). However, for large scale arrays, targets are more likely to be located in the near-field region, and the USW assumption should be used [50, 51], as shown in Fig. 5(b). Let the location of target $k$ be represented as ${{\mathbf{q}}_{k}}={[r_{k}\cos\theta_{k},r_{k}\sin\theta_{k}]^{T}},k=1,...,K$, where $r_{k}$ is the distance between target $k$ and a given reference antenna, and $\theta_{k}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ is the angle of target $k$ with respect to the normal vector of the array. Then, the receive steering vector $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ can be expressed as
$\left[{\mathbf{a}}_{\mathrm{R}}({{\mathbf{q}}_{k}})\right]_{m}={{e^{j\frac{{2\pi}}{\lambda}\left({{r_{k,m}}-{r_{k}}}\right)}}},m=1,...,M_{r},$ (10)
where $r_{k,m}$ denotes the distance between the $k$-th target and the $m$-th receive antenna. For narrowband signals, half-wavelength inter-antenna spacing is typically assumed, i.e., $d=\frac{\lambda}{2}$. If we assume the $r$-th antenna is the reference located at the origin of the array’s coordinate system, the positions of the $M_{r}$ antennas can be written as $[0,\varepsilon_{m}d],m=1,\cdots,M_{r}$, where $|\varepsilon_{m}d|$ is the distance between the $m$-th antenna and the reference, and $\varepsilon_{r}=0$. Thus, $r_{k,m}$ can be written as
@@
${{\mathbf{a}}_{{\mathrm{R}}}}({\theta_{k}})=\left[{{e^{-j\frac{{2\pi}}{\lambda}\varepsilon_{m}d\sin{\theta_{k}}}}}\right],m=1,...,M_{r}.$ (13)
@@
Comparing (13) and (14), we see that in the far-field, only the angle can be estimated from one snapshot signal on different antenna elements, while both target angle and distance can be estimated in the near-field. The general expression in (14) can be simplified using a second-order Taylor approximation of (11) [51], leading to
@@
$\;{\bf{a}}_{\rm{R}}({r_{k}},{\theta_{k}})=\left[{{e^{j\left({{\omega_{k}}{\varepsilon_{m}}+{\psi_{k}}\varepsilon_{m}^{2}}\right)}}}\right],m=1,...,M_{r},$ (16)
@@
In Sections IV and V, we will present typical parameter estimation algorithms based on the four methods presented in Fig. 4, first for far-field and then for near-field scenarios.

### v2

@@ III-A Unified ISAC Model
  • UE-to-BS bistatic ISAC mode: For UE-to-BS bistatic ISAC mode, the ISAC-TX in Fig. 3 corresponds to a UE, while the Com-RX and Sen-RX are collocated at the BS, as shown in Fig. 4(b). Note that unlike the BS monostatic mode in Fig. 4(a), since only signal reception needs to be performed at the BS in Fig. 4(b), the Com-RX and Sen-RX may use the same set of antennas. However, synchronization is essential due to the use of different local oscillators (LO) and reference clocks by the ISAC BS and the UE. The ISAC signal $\mathbf{x}(t)$ in (1)-(3) is transmitted by the UE, and is denoted as $\mathbf{x}_{\mathrm{UE}}(t)\in\mathbb{C}^{M_{t}\times 1}$.
@@
  • BS-to-BS bistatic ISAC mode: For BS-to-BS bistatic ISAC mode, the ISAC-TX in Fig. 3 corresponds to a BS, while the Com-RX and Sen-RX occur at a UE and another BS, respectively, as shown in Fig. 4(c). Note that unlike the UE-to-BS bistatic mode in Fig. 4(b), although the BS bistatic mode also requires synchronization, it can be achieved through a wired connection between the two ISAC BSs. Furthermore, the ISAC signal $\mathbf{x}(t)$ in (1)-(3) is broadcast by the transmit BS, and is denoted as $\mathbf{x}_{\mathrm{BS}}(t)\in\mathbb{C}^{M_{t}\times 1}$.
Figure 5: Primary signal processing operations for monostatic and bistatic MIMO-OFDM ISAC systems.
@@
The corresponding ISAC-TX, Com-RX, and Sen-RX of the unified model in Fig. 3, as well as the required processing for the six ISAC modes, are summarized in Table IV.
@@
In the following, we focus on ISAC systems implemented with OFDM waveforms, which is the dominant waveform in 4G and 5G communication systems, and will continue to play the same role for 6G. In OFDM-ISAC systems, target sensing can be performed based on pilots only or on the fully reconstructed transmitted OFDM data frame. The former approach can be easily implemented, but its sensing performance is limited by the availability of time-frequency resources allocated to pilots. The latter can fully exploit all of the time-frequency resources for target sensing using known (for monostatic sensing) or demodulated communication symbols (for bistatic sensing), at the expense of increased complexity. In this paper, we employ the OFDM frame-based approach for sensing tasks using all time-frequency resources allocated to both pilots and communication data.
@@ III-B MIMO-OFDM ISAC Signal Processing and Modeling
For MIMO-OFDM ISAC, the ISAC-TX transmits OFDM signals over $N$ subcarriers and $P$ OFDM symbols. The subcarrier spacing and the OFDM symbol duration without the CP are denoted by $\Delta f$ and $T$, respectively, where $T=\frac{1}{\Delta f}$. The system bandwidth is thus $B=N\Delta f$. Let $T_{\mathrm{cp}}$ denote the CP duration, so the OFDM symbol duration including CP is $T_{s}=T+T_{\mathrm{cp}}$. To avoid ISI for communication, the CP duration should be no smaller than the multi-path delay spread $\tau_{\mathrm{com}}$ of the communication channel $\mathbf{H}_{c}(t,\tau)$, since the Com-RX only begins receiving communication signals upon receiving its first path. However, for target sensing, the CP duration should be larger than the propagation delay $\tau_{\mathrm{sen}}$ corresponding to the ISI-free sensing distance [102], which is different from the communications case since the Sen-RX begins to receive signals as soon as the ISAC-TX transmits, so that no nearby targets are missed. Thus, in ISAC systems, the CP should be set as $T_{\mathrm{cp}}\geq\tau_{\max}$, where $\tau_{\max}=\max\{\tau_{\mathrm{com}},\tau_{\mathrm{sen}}\}$. The OFDM signal $\mathbf{x}(t)\in\mathbb{C}^{M_{t}\times 1}$ transmitted by the $M_{t}$-antenna ISAC-TX array is given by
@@
$\qquad\quad\times e^{j2\pi\upsilon_{k}t}\mathrm{rect}\left(\frac{t-\tau_{k}-pT_{s}}{T_{s}}\right)+\mathbf{z}(t),$ (5)
@@
The Sen-RX samples the sensing signal with sampling rate $B=N\Delta f$. To avoid inter-carrier interference (ICI) caused by large Doppler frequency shifts, the subcarrier spacing is typically assumed to be much larger than the Doppler frequency shift, i.e., $\Delta f\gg|\upsilon_{k}|,\forall k=1,2,\cdots,K$. Under this assumption, the $q$-th ($q=0,1,\cdots,N-1$) sample of the $p$-th block can be expressed as
@@
$\qquad\qquad\times e^{j2\pi\upsilon_{k}\left(q/B+pT_{s}+T_{\mathrm{cp}}\right)}+\mathbf{z}^{p}\left[q\right]$
@@
$\qquad\qquad\qquad\qquad\times e^{j2\pi pT_{s}\upsilon_{k}}+\mathbf{z}^{p}[q],$
@@
The main signal processing operations associated with MIMO-OFDM ISAC systems for the monostatic and bistatic modes are illustrated in Fig. 5. The operations for the monostatic and bistatic ISAC modes mainly differ in three aspects. First, in the monostatic mode, the ISAC-TX and Sen-RX use the same LO and reference clock, avoiding the need for synchronization with no STO or CFO in this case. By contrast, the bistatic mode requires different LOs and reference clocks, necessitating synchronization at the RX. Second, unlike the bistatic case, the monostatic ISAC-TX and Sen-RX operate in full-duplex mode and require SI cancellation. Finally, the information-bearing symbols transmitted by the ISAC-TX are available at the monostatic Sen-RX since they are collocated. However, in the bistatic case, only pilot data is available, and the information-bearing symbols need to be demodulated before target sensing. Synchronization between basestations in bistatic systems can be achieved via both wireless and wired connections, and fine synchronization can be achieved after FFT demodulation to further reduce the residual STO and CFO in scenarios requiring higher accuracy.
@@
Parameter estimation is one of the most important sensing tasks in ISAC systems, as it underpins functions like target localization, tracking, and imaging. Therefore, in the following, we will focus on estimation of the location $\mathbf{q}_{k}$ (or the angle $\theta_{k}$ in the far-field case), the propagation delay $\tau_{k}$, and the Doppler frequency shift $\upsilon_{k}$ of the sensing targets. Generally speaking, the target parameters are estimated from the spatial, delay, and Doppler domain data. Inspired by traditional joint parameter estimation methods [103, 104], we propose four frameworks to estimate target parameters across the three domains. As shown in Fig. 6, depending on whether the parameters are estimated jointly or separately in each domain, the four frameworks are parallel one-domain, sequential one-domain, joint two-domain, and joint three-domain parameter estimation. The parallel one-domain approach separately estimates the sets of target locations $\mathbb{S}_{\mathbf{q}}=\{\mathbf{q}_{1},\cdots,\mathbf{q}_{K}\}$ or angles $\mathbb{S}_{\theta}=\{{\theta}_{1},\cdots,{\theta}_{K}\}$, delays $\mathbb{S}_{\tau}=\{\tau_{1},\cdots,\tau_{K}\}$ and Doppler shifts $\mathbb{S}_{\upsilon}=\{\upsilon_{1},\cdots,\upsilon_{K}\}$, and then associates the estimated parameters into groups $(\mathbf{q}_{k}/\theta_{k},\tau_{k},\upsilon_{k})$ for each of the $K$ sensing targets. This approach has a relatively low computational complexity, but its sensing performance is limited by potential errors in the parameter grouping.
@@
On the other hand, the sequential one-domain approach performs spatial, delay, and Doppler domain estimation sequentially. For example, spatial domain estimation can be applied first to determine the $K_{\hat{\mathbf{q}}}$ distinguishable locations/angles. When the spatial resolution is limited, it may not be possible to distinguish all $K$ targets in the spatial domain alone. Thus, in general, $K_{\hat{\mathbf{q}}}\leq K$. After location/angle estimation, beamforming can be applied to separate the $K_{\hat{\mathbf{q}}}$ signals in the spatial domain, denoted as $\mathbf{X}_{k_{\hat{\mathbf{q}}}},k_{\hat{\mathbf{q}}}=1,\cdots,K_{\hat{\mathbf{q}}}$. Subsequently, delay and Doppler domain processing can be applied. Note that the order in which the different types of parameters are estimated depends on the resolution and target separation in the three domains. The sequential one-domain estimation method does not require parameter grouping, but its complexity is greater than parallel one-domain processing.
@@
The details of each method will be discussed in Sections IV and V for the far-field and near-field cases, respectively. Note that the TX and RX steering vectors $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})$ and $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ are functions of target location $\mathbf{q}_{k}$, which in general depends on both the angle of departure (AoD)/AoA $\theta_{k}$ and distance $r_{k}$ between the target and the reference antenna. In the following, we present expressions for the RX steering vectors for far-field and near-field scenarios. Similar results can be obtained for the TX array response.
@@ III-C Array Response Vector: From Far-Field to Near-Field
When the BS array aperture is small, targets are typically located in the far-field region, and the conventional UPW model accurately characterizes the variations of signal phases across the array, as shown in Fig. 7(a). However, for large-scale arrays, targets are more likely to be located in the near-field region, and the USW assumption should be used [50, 51], as shown in Fig. 7(b).
@@
(b) Uniform spherical wave.
@@
Let the location of target $k$ be represented as ${{\mathbf{q}}_{k}}={[r_{k}\cos\theta_{k},r_{k}\sin\theta_{k}]^{T}},k=1,...,K$, where $r_{k}$ is the distance between target $k$ and a given reference antenna, and $\theta_{k}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ is the angle of target $k$ with respect to the normal vector of the array. Then, the receive steering vector $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ can be expressed as
$\left[{\mathbf{a}}_{\mathrm{R}}({{\mathbf{q}}_{k}})\right]_{m}={{e^{j\frac{{2\pi}}{\lambda}\left({{r_{k,m}}-{r_{k}}}\right)}}},m=1,.,M_{r},$ (10)
where $r_{k,m}$ denotes the distance between the $k$-th target and the $m$-th receive antenna. For narrowband signals, half-wavelength inter-antenna spacing is typically assumed, i.e., $d=\frac{\lambda}{2}$. If we assume the $r$-th antenna is the reference located at the origin of the array’s coordinate system, the positions of the $M_{r}$ antennas can be written as $0,\varepsilon_{m}d],m=1,\cdots,M_{r}$, where $|\varepsilon_{m}d|$ is the distance between the $m$-th antenna and the reference, and $\varepsilon_{r}=0$. Thus, $r_{k,m}$ can be written as
@@
${{\mathbf{a}}_{{\mathrm{R}}}}({\theta_{k}})=\left[{{e^{-j\frac{{2\pi}}{\lambda}\varepsilon_{m}d\sin{\theta_{k}}}}}\right],m=1,.,M_{r}.$ (13)
@@
Comparing (13) and (14), we see that in the far-field, only the angle can be determined from one snapshot on different antenna elements, while both target angle and distance can be estimated in the near-field. The general expression in (14) can be simplified using a second-order Taylor approximation of (11) [51], leading to
@@
$\;{\bf{a}}_{\rm{R}}({r_{k}},{\theta_{k}})=\left[{{e^{j\left({{\omega_{k}}{\varepsilon_{m}}+{\psi_{k}}\varepsilon_{m}^{2}}\right)}}}\right],m=1,.,M_{r},$ (16)
@@
In Sections IV and V, we will present typical parameter estimation algorithms applicable to the four methods presented in Fig. 6, first for far-field and then for near-field scenarios.
III-D Prototype Design and Hardware Requirements
In this section, we first introduce a typical MIMO-OFDM ISAC prototype system. Then, we elaborate on the fundamental hardware requirements and key parameters of ISAC prototype designs, including servers, development software, field programmable gate array (FPGA) and radio frequency (RF) development kits, and MIMO array design.
III-D1 Monostatic MIMO-OFDM ISAC Prototype
The authors in [107] describe a mmWave MIMO-OFDM monostatic ISAC prototype that they designed for internet of vehicles (IoV) applications. The ISAC BS provides downlink multimedia data to vehicles, while simultaneously utilizing the OFDM echoes to sense the position and velocity of the vehicles and surrounding obstacles to provide obstacle avoidance services. As shown in Fig. 8, the ISAC BS is equipped with one universal software radio peripheral (USRP) and two mmWave phased arrays (PSAs) that function as ISAC-TX and Sen-RX, respectively. The vehicle is equipped with one USRP and a PSA, serving as Com-RX. At the ISAC BS, the USRP is responsible for generating OFDM ISAC signals, beam and vehicle control signaling, and sensing signal processing, while the two PSAs separately handle signal transmission and reception. On the vehicle side, the PSA receives communication signals, and the USRP performs communication signal processing, demodulates the video data stream and control commands, and forwards them to the vehicle systems. In the following, we discuss the fundamental hardware requirements and critical parameters for ISAC prototype development based on this system.
Figure 8: Monostatic MIMO-OFDM ISAC prototype system architecture [107].
Figure 9: Hardware architecture of ISAC transceiver terminal.
III-D2 Hardware Requirements and Key Parameters
The basic architecture of the BS and UE prototypes is illustrated in Fig. 9. Depending on their distinct roles in different ISAC modes, such as ISAC-TX, Com-RX, and Sen-RX, their transceiver arrays can operate either simultaneously or separately. For instance, as depicted in Fig. 8, the transmit and receive arrays of the ISAC BS operate concurrently, while only the receive array is active at the UE. Next, we introduce the hardware components and their key parameters.
Servers: Servers are employed to fulfill data storage requirements, computing power support, software configuration and updates. The hardware configuration of the host, including the central processing unit (CPU), random access memory (RAM), and solid state drive (SSD), significantly impacts the operational efficiency, data processing, and storage capabilities of the experimental system. High-performance PCs generally feature a relatively large physical size, rendering them unsuitable for deployment on compact devices such as drones or mobile terminals. An effective solution is to deploy cloud servers for computing and storage, thereby enhancing computing power and storage resources. However, the latency induced by data upload and result retrieval may compromise the real-time performance of the system.
Development software: Development software comprises various programming tools designed to drive hardware operations, enabling core functionalities such as system parameter configuration, ISAC waveform parameter setup, C&S signal processing, human-machine interaction, and result visualization. Specifically, the Verilog programming language facilitates the implementation of signal processing functionalities on FPGA chips, enabling high-efficiency and low-power processing. This approach enhances both the efficiency and real-time performance of the experimental system, albeit at the cost of high workload and implementation complexity. LabVIEW and GNU Radio are two graphical programming development environments that incorporate extensive libraries of functional modules tailored for communication and radar signal processing, effectively reducing the development challenges of experimental systems. LabVIEW is a commercial software platform developed by National Instruments (NI), while GNU Radio is an open-source software framework. In the prototype system illustrated in Fig. 8, the authors implemented the core OFDM signal processing flow shown in Fig. 5 via the FPGAs, while more complex sensing signal processing was handled by LabVIEW running on the server. This strategy effectively balances system efficiency with development complexity.
FPGA and RF development kits: The transceiver hardware primarily comprises FPGA chips for signal and data processing, along with various RF modules such as digital-to-analog converters (DAC), analog-to-digital converters (ADC), filters, amplifiers, and up/down converters. For prototype system development, commercial software-defined radio (SDR) platforms can be used, such as the USRP product series from NI. Several USRP hardware parameters are particularly critical for ISAC performance, including carrier frequency range, maximum bandwidth, and the number of TX/RX RF chains. Most current USRP products cover the sub-6 GHz frequency band, while experimental validation for higher frequency bands, such as mmWave and THz, requires additional high-frequency up/down converters. The system bandwidth supported by USRP restricts the communication rate and delay resolution of ISAC prototype systems. Finally, the number of USRP TX/RX RF chains is generally limited. For experimental systems requiring more RF chains, such as massive MIMO or XL-MIMO ISAC systems, multiple USRPs need to be jointly utilized with data aggregation and synchronization. Additionally, the server must be equipped with the USRP hardware driver (UHD) to facilitate data transmission and control signaling interaction with the USRP via interfaces such as USB, Ethernet, or PCIe.
MIMO array design: The antenna array is another critical hardware component in MIMO-OFDM ISAC systems, determining the near-field sensing region and spatial sensing resolution. Depending on the implementation methods used, MIMO arrays can be categorized as digital MIMO arrays, phased MIMO arrays, movable antenna arrays [60], switched arrays, and ray antenna arrays (RAAs) [108, 109]. For a digital MIMO array, each antenna is connected to an individual RF chain, enabling precise digital beamforming at both transmitter and receiver. It supports diversity transmission at the transmitter and spatial sensing through array signal processing at the receiver. The theory described in this paper is based on the assumption that digital MIMO arrays are deployed. For large-scale or extremely large-scale arrays, the number of required RF chains increases correspondingly, leading to a sharp rise in hardware costs. Phased MIMO arrays significantly reduce the number of required RF chains, down to a minimum of only one RF chain, but require a dedicated phase shifter for each antenna. Phased MIMO arrays usually rely on hybrid digital-analog or analog beamforming for signal transmission and reception, and beam scanning replaces the spatial signal processing shown in Fig. 6. The analog beamforming precision is limited by the phase shifters, resulting in degraded C&S performance compared to digital MIMO arrays, albeit at a lower hardware cost. Movable antenna arrays and switched arrays are designed based on the principles of synthetic aperture radar (SAR) [110]. Movable antenna arrays rely on physical movement of antennas, while switched arrays use RF switches to achieve high-speed switching between array elements. Both approaches can achieve extremely large-scale aperture sensing at very low cost. However, because such architectures require time-division signal transmission and reception, they are only suitable for low-mobility scenarios and exhibit limited Doppler resolution and beamforming gain. The RAA is composed of multiple simple uniform linear arrays (ULAs) whose elements are connected to form a fixed beam at the array boresight without adaptive digital/analog beamforming, thus significantly reducing the hardware cost. Through careful design of the ULA orientations, this approach can achieve uniform angular resolution and enable the use of antenna elements with higher directionality, thus demonstrating considerable potential for C&S applications. However, potential shadowing and mutual coupling effects among the ULAs may arise, necessitating specialized designs to mitigate these issues. Furthermore, antennas and arrays in the sub-6G band are technologically mature and cost-effective, whereas higher-frequency antennas and arrays are considerably more expensive. Consequently, some research efforts are currently focused on developing low-cost and low-power high-frequency antenna arrays, such as holographic radar arrays [111, 112, 113] and reconfigurable antenna arrays [114, 115].
III-E Lessons Learned
The model depicted in Fig. 3 provides a unified representation for the six commonly used ISAC modes illustrated in Fig. 2. This unified model helps simplify the C&S signal model derivation and signal processing workflow design, while also offering a foundational framework for ISAC system design and prototype validation. However, different ISAC modes have distinct channel characteristics. For instance, compared to UE-side sensing, an ISAC BS is more likely to be equipped with massive or XL-MIMO arrays. Consequently, targets in BS-side sensing are more likely located in the near-field region, necessitating the consideration of spherical wave propagation models and near-field spatial signal processing. In terms of coverage, monostatic sensing typically exhibits circular or spherical coverage patterns, whereas bistatic sensing generally follows Cassini oval curves or surfaces [46, 116]. Finally, for the same target, the RCS may differ between monostatic and bistatic sensing due to variations in target scattering characteristics. Thus, single-mode ISAC may lead to detection blind spots or poor sensing performance, whereas multi-mode collaborative ISAC can enhance target detection accuracy and parameter estimation performance [117].
For prototype development, beyond the fundamental hardware requirements discussed in Section III-D, it is essential to design the ISAC system architecture as shown in Fig. 8 based on application scenarios and functional requirements. This involves configuring hardware parameters and time-frequency resources, designing signal processing workflows, and feedback mechanisms for upper-layer applications. The authors in [118] proposed a five-layer architecture to address the requirements and functions of ISAC systems at various levels, encompassing the terminal and signal layer, resource layer, function layer, application layer, and computation layer, and they conducted survey on existing ISAC prototype validation efforts. In general, significant progress has been made in the development of monostatic[119, 107, 25], bistatic [37], and cell-free multistatic [120] MIMO-OFDM ISAC prototypes, as well as in OFDM-based simultaneous localization and mapping (SLAM) [9], SAR imaging [121], and ISAC-enabled CKM [32] systems. However, most of these efforts focus on the far-field case, while near-field scenarios involving massive MIMO or XL-MIMO require further research into corresponding prototype designs and experimental validations.